OpenStudy (anonymous):

Write the following numbers in standard notation 3.102 x 102 = 1.75 x 10-3 = Write the following numbers in scientific notation 120,000 = 12 = Write 0.000001825 in scientific notation= Write 4.9 x 102 in standard form=

OpenStudy (anonymous):

3.102[10^{2}]=310.2 1.75*[10^{-3}]=0.00175 120000=1.2*10^5 0.000001825=1.825*10^-6

OpenStudy (anonymous):

thank u!